# TTPXHunter: Actionable Threat Intelligence Extraction as TTPs form Finished Cyber Threat Reports

### 05 Mar 2024 v1

@@ IV-A Contextual Data Augmentation
We can see at step 1⃝ in Fig. 2, an example of input sentence as "Adversary obtained credentials using compromised systems". In this case, if we mask the word "Adversary", then the model predicts possible words that preserve the contextual meaning (step 2⃝ in Fig. 2). The probabilities associated with each word indicate the confidence level of the predicted word. We select the top $5$ words from this list and generate five new sentences for each input sentence (step 3⃝ in Fig. 2). Now, there is a chance that the sentence’s meaning may deviate after replacing the masked word with the predicted word. Therefore, we employ cosine similarity to compare each newly generated sentence with the original sentence (step 4⃝ in Fig. 2) and select the sentence with the highest similarity index. To compute the cosine similarity, we generate contextual embedding for both the original and generated sentences using Sentence Transformer [11]. The similarity score ranges between 00 (not similar) to $1$ (Exactly similar), and we set a threshold $(\theta)$ of $0.975$. We retain only those sentences having similarity scores greater than $\theta$. The sentence augmentation algorithm is outlined in Algorithm 1, which we employ to construct the augmented sentence-TTP dataset. More detailed information regarding the distribution of the augmented sentence-TTP dataset is present in the section V-A.
@@ IV-B Prepossessing & Fine-Tuning
The finetuning process is similar to the TTPHunter’s finetuning. As TTPHunter extracts sentence embedding from the traditional BERT model and passes it to a linear classifier for classification, we employ a similar method in TTPXHunter. It first extracts domain-specific embedding using the SecureBERT language model and passes embedding to a linear classifier for TTP classification. The finetuning process is meticulously orchestrated to optimize performance on the processed sentence dataset. We initiate the process by configuring essential parameters, setting the learning rate to $52e-$ 5 to ensure subtle adjustments to the model, choosing a batch size of $64$ to balance computational efficiency with training effectiveness, and limiting the training to $10$ epochs to prevent overfitting. We leverage the Hugging Face’s Transformers library [8] for its robust support of transformer models. Sentences are tokenized using SecureBERT’s tokenizer, aligning with its pre-trained understanding of language structure, and inputs are standardized to a maximum length of $256$ tokens. We finetune our model to let it understand the context for ATT&CK TTP’s leverage the understanding in performing classification. The model’s finetuning is shown in Fig. 3.
@@ IV-C TTP Extraction
TTPXHunter takes finished threat reports as input and breaks the report contents into a list of individual sentences. Then, each sentence is processed and passed to the TTPXHunter’s finetuning architecture for classification. The system classifies each sentence to identify TTPs. Finally, it aggregates these TTPs and creates a list of TTPs extracted from the analyzed report. A threat report usually contains many irrelevant sentences that do not reflect TTPs. However, our classification module is a closed-world solution (the input will surely be classified into at least one target class). As a result, irrelevant sentences can also be classified as one of the TTPs in the target class, which can lead to huge false positives. To reduce this effect, we filter irrelevant sentences, which is similar to TTPHunter. We employ a threshold mechanism to filter irrelevant sentences. The classification module only maps the TTP class for a given sentence representing any TTP, i.e., a relevant sentence. We filter irrelevant sentences based on classification confidence score. We fix a threshold $(\Theta)$ and filter the sentences if the classifier’s confidence score is below the threshold. The classifier only maps to the TTP class if the confidence score is greater than the fixed threshold $(\Theta)$. We follow the same threshold value experimentally obtained in TTPHunter, i.e., $0.644$.
Once the linear classifier extracts the list of TTPs present in the threat report, TTPXHunter converts the list to the Structured Threat Information eXpression (STIX) [15] format, significantly enhances cyber threat intelligence operations. It supports detailed analysis and correlation of threats by providing a rich, structured representation of cyber threat information. This structured format facilitates automated processing and threat response, increasing operational efficiency. Moreover, the consistency and standardization offered by STIX improve communication within and between organizations, ensuring a common understanding of cyber threats.
@@ V Experiments and Results
This section presents the details about the prepared dataset and the experiments. We also discuss the results obtained and chosen performance measures.
@@ V-A Dataset
We compute model performance on our test dataset in the form of sentence-to-TTP mapping. However, in the real world, the threat data is in the form of reports, containing a set of sentences mixed with relevant and irrelevant sentences. Therefore, we also evaluate the model’s performance report-wise. As a result, we prepared two datasets to measure the efficiency of TTPXHunter and the current literature: The augmented Sentence-TTP dataset and the Report-TTP dataset.
@@ V-A1 Augmented Sentence-TTP Dataset
To prepare the sentence-TTP dataset, we consider the TTPHunter dataset as a base dataset, prepared using MITRE ATT&CK knowledgebase [1]. The base dataset consists of two columns: sentences and their corresponding TTP ID. The dataset consists of $10,906$ sentences over $193$ TTPs. We use our proposed data augmentation algorithm, explained in Algorithm 1, and prepare the dataset consists $39,296$ sentences over $192$ TTP classes. The TTP ID $T1059$ (Command and Scripting Interpreter) consists of the highest number of sentences as $800$, and TTP ID $T1127$ (Trusted Developer Utilities Proxy Execution) consists of the lowest number of samples as $3$. On average, the number of samples in our dataset is $203$. The distribution of data samples is present in Appendix VIII and a glimpse of the dataset is shown at step 5⃝ in Fig 2.
@@ V-A2 Report-TTP Dataset
Generally, threat data is present in the form of threat analysis reports. We evaluate TTPXHunter on a document dataset, which demonstrates the performance of TTPXHunter in filtering irrelevant sentences from threat reports. It also tells us how efficiently TTPXHunter can extract threat intelligence from threat reports. We manually collect $149$ threat reports published by various prominent security firms. We manually label a set of TTPs present in each report to create a ground truth. The document-TTP dataset contains two columns: threat report and list TTPs in the corresponding threat report.
@@ V-B Evaluation
We evaluate TTPXHunter using various performance metrics and compare its performance with several current studies. As our dataset is imbalanced, shown in Fig. 6, we consider macro-averaged precision, recall, and f1-score. This approach ensures balanced evaluation across all classes. By giving equal importance to each TTP class, these metrics prevent the majority class’s dominance from overshadowing the minority class’s performance. This approach promotes the development of effective and fair models across different types of TTP sentences and is essential for nuanced TTP classification tasks.
@@ V-B1 Augmented Sentence-based Evaluation
We finetune TTPXHunter on the prepared augmented dataset and compare the other two BERT-based models, i.e., TRAM [7] and TTPHunter [1] present in the literature. We divide the prepared augmented dataset in train and test set in $80:20$ ratio.
TTPXHunter vs TRAM : We fine tune the TTPXHunter on the train set and evaluate its performance using the chosen performance metrics. Further, we finetune the literature TRAM [7] on the same dataset and evaluate its performance. By employing TRAM on the augmented dataset, we extend the capability of TRAM from the 50 most frequently used TTPs to the full spectrum of TTPs. As a result, it gives common ground for comparing TTPXHunter and TRAM. The result obtained by both methods and their comparison is shown in Fig. 4. As we can see, the TTPXHunter outperforms the TRAM, which reflects the difference between contextual embedding of general scientific BERT and domain-specific BERT embedding capability. This result reflects that the domain-specific language model provides a better contextual understanding of embedding than the language model trained on general scientific terms.
@@
TTPXHunter vs TTPHunter : We assess the performance of our proposed TTPXHunter alongside state-of-the-art TTPHunter. Our proposed extraction method, TTPXHunter, performs better than state-of-the-art TTPHunter. TTPXHunter’s superiority is due to using a cyber-domain-specific finetuned language model. Sentences containing domain-specific terms, such as ”window” and ”registry,” introduce a distinct context that differs from general English. This distinction allows our method, built upon a domain-specific language model, to capture and interpret the contextual meaning more accurately than traditional models. In addition, TTPXHunter can identify the range of $193$ TTPs with $0.92$ f1-score, whereas TTPHunter is limited to only $50$ TTPs. The improvement in the result and the capability to identify all ranges of TTPs make TTPXHunter superior to TTPHunter.
@@
Further, to understand the efficiency of the data augmentation method and compare the performance of TTPXHunter and TTPHunter on the same base, we evaluate TTPXHunter on the ground of TTPHunter. We only selected 50 TTP sets for which we developed TTPHunter and evaluated both. The obtained result is shown in Fig. 5. As we can see, TTPXHunter outperforms TTPHunter on the 50-TTP set ground of TTPHunter. It reflects that augmenting more samples for a 50-TTP set and employing a domain-specific language model enables the classifier to understand the context of TTP better.
@@ V-B2 Report-based Evaluation
In the real world, we have threat reports in the form of natural language rather than sentence-wise datasets, and these reports contain information along with TTP-related sentences. So, we evaluate the TTPXHunter on the report dataset, which contains each sample as threat report sentences and a list of TTP explained in the report. Extracting TTPs from threat reports is a multi-label problem because for any given sample, i.e., threat report in this case, a list of TTP classes is expected as output. The evaluation of such classification also requires careful consideration because of multi-label classification.
Evaluation Metrics: In the multi-label problem, the prediction vector appears as a multi-hot vector rather than a one-hot vector in a multi-class problem. In the multi-label case, there can be a situation where not all expected TTP classes were predicted; instead, a subset of them is correctly predicted. However, as the whole multi-hot vector does not match, the prediction is considered wrong. For example, If the true label set contains $\{T1,T2,T3\}$ and the predicted label is $\{T2,T3\}$, then it can be considered false even if $T2$ and $T3$ are correctly classified, as all true value has not been classified. So, relying on accuracy may not be a good choice for multi-label problems [22]; instead, we consider hamming loss as a performance metric to deal with such a scenario. The hamming loss measures the error rate label-wise [16, 17, 22]. It calculates the ratio of incorrect labels to all labels. For given $k$ threat reports, the hamming loss is defined as:
@@
Further, we also evaluate macro precision, recall, and f1-score by leveraging a multi-label confusion matrix package from sklearn [10]. Then, we calculate true positive, false positive, and false negative for each class and calculate these performance metrics for each class. Further, we calculate the macro average between all classes to get macro-averaged performance metrics for all chosen measures, i.e., precision, recall, and f1-score. We prefer the macro-average method to ignore biases towards the majority class and provide equal weight overall classes.
We consider four state-of-the-art methods, i.e., [9, 5, 3, 7] for the comparison against TTPXHunter based on these metrics over the report dataset. This comparison aims to understand the effectiveness of TTPXHunter over state-of-the-art for TTP extraction from finished threat reports. These methods provide the list of TTPs extracted from the given threat report and the model confidence score for each. Out of all extracted TTPs, only relevant TTPs are selected based on the threshold mechanism decided by each method. We evaluate the state-of-the-art method’s performance based on the threshold value given in their respective paper. For TTPXHunter, we obtain the same threshold we experimentally chose for TTPHunter, i.e., $0.644$. The result obtained from all implemented methods on our report dataset is shown in Table IV. We can see that the presented method, TTPXHunter, outperforms all implemented methods in every chosen metric. The TTPXHunter achieves the lowest hamming loss and the highest other performance metrics in our evaluation. This performance superiority over state-of-the-art methods demonstrates the efficiency of the TTPXHunter, and we plan to make it open for the benefit of the community.
@@ VI Limitations & Future Directions
In addition to the advantages of TTPXHunter, like threat intelligence extraction, threat profiling, and sharing, it has some limitations that one should be careful about while using it for actionable insights. The MITRE ATT&CK framework is not a kind of fixed knowledge base. Instead, MITRE threat researchers are continuously updating it. One may need to finetune the model again if new TTPs are added to get the newer TTPs as predicted. While finetuning on newer TTPs may require using our proposed method to create new augmented sentences specific to newer TTPs. Therefore, we plan to make it public so one can adapt our method for any new upcoming versions of the ATT&CK framework.
Further, the TTPXHunter contains a one-to-one classifier model, which maps a given sentence to a single TTP. Like a single sentence, a sentence containing one-to-many mapping may explain more than one TTP. For example, the sentence is ”The attacker gained initial access through a phishing email and obtained persistence via run registry modification.” TTPXHunter maps this sentence to T1566 (Phishing) or T1037 (Boot or Logon Initialization Scripts) in such a scenario. Extending the capability of TTPXHunter to identify such one-to-many mapping can help us improve the model’s performance. We plan to take up this challenge to improve the efficiency of TTPXHunter in the future.

### 21 Mar 2024 v3

@@ IV-A Contextual Data Augmentation
We can see at step 1⃝ in Fig. 2, an example of input sentence as "Adversary obtained credentials using compromised systems". In this case, if we mask the word "Adversary", then the model predicts possible words that preserve the contextual meaning (step 2⃝ in Fig. 2). The probabilities associated with each word indicate the confidence level of the predicted word. We select the top $5$ words from this list and generate five new sentences for each input sentence (step 3⃝ in Fig. 2). The sentence’s meaning may deviate after replacing the masked word with the predicted word. Therefore, we employ cosine similarity to compare each newly generated sentence with the original sentence (step 4⃝ in Fig. 2) and select the sentence with the highest similarity index. To compute the cosine similarity, we generate contextual embedding for both the original and generated sentences using Sentence Transformer [13]. The similarity score ranges between 00 (not similar) to $1$ (Exactly similar), and we set a threshold $(\theta)$ of $0.975$. We retain only those sentences having similarity scores greater than $\theta$. The sentence augmentation algorithm is outlined in Algorithm 1, which we employ to construct the augmented sentence-TTP dataset. More detailed information regarding the distribution of the augmented sentence-TTP dataset is present in the section V-A.
@@ IV-B Prepossessing & Fine-Tuning
The finetuning process is similar to the TTPHunter’s finetuning. As TTPHunter extracts sentence embedding from the traditional BERT model and passes it to a linear classifier for classification, we employ a similar method in TTPXHunter. It first extracts domain-specific embedding using the SecureBERT language model and passes embedding to a linear classifier for TTP classification. The finetuning process is meticulously orchestrated to optimize performance on the processed sentence dataset. We initiate the process by configuring essential parameters, setting the learning rate to $51e-$ 5 to ensure subtle adjustments to the model, choosing a batch size of $64$ to balance computational efficiency with training effectiveness, and training it for $10$ epochs. We leverage the Hugging Face’s Transformers library [9] for its robust support of transformer models. Sentences are tokenized using SecureBERT’s tokenizer, aligning with its pre-trained understanding of language structure, and inputs are standardized to a maximum length of $256$ tokens. We finetune our model to let it understand the context for ATT&CK TTP, which further enhances the classification capability. The model’s finetuning is shown in Fig. 3.
@@ IV-C TTP Extraction
TTPXHunter takes finished threat reports as input and breaks the report contents into a list of individual sentences. Then, each sentence is processed and passed to the TTPXHunter’s finetuning architecture for classification. The system classifies each sentence to identify TTPs. Finally, it aggregates these TTPs and creates a list of TTPs extracted from the analyzed report. A threat report usually contains many irrelevant sentences that do not reflect TTPs. However, our classification module is a closed-world solution (the input will surely be classified into at least one target class). As a result, irrelevant sentences can also be classified as one of the TTPs in the target class, which can lead to huge false positives. To reduce this effect, we filter irrelevant sentences, similar to TTPHunter. We employ a threshold mechanism in the classification module to filter irrelevant sentences and only map sentences that explain TTPs, i.e., a relevant sentence. We filter irrelevant sentences based on classification confidence score. We fix a threshold $(\Theta)$ and filter the sentences if the classifier’s confidence score is below the threshold. We follow the same threshold value experimentally obtained in TTPHunter, i.e., $0.644$.
Once the linear classifier extracts the list of TTPs present in the threat report, TTPXHunter converts the list to the Structured Threat Information eXpression (STIX) [18] format, significantly enhances cyber threat intelligence operations. It supports detailed analysis and correlation of threats by providing a rich, structured representation of cyber threat information. This structured format facilitates automated processing and threat response, which increases operational efficiency. Moreover, the consistency and standardization offered by STIX improve communication within and between organizations to ensure a common understanding of cyber threats.
@@ V Experiments and Results
This section presents details about the prepared dataset and the experiments. We also discuss the results obtained and chosen performance measures.
@@ V-A Dataset
We compute model performance on our test dataset through sentence-to-TTP mapping. However, in the real world, the threat data is in the form of reports containing a set of sentences mixed with relevant and irrelevant sentences. Therefore, we also evaluate the model’s performance report-wise. As a result, we prepared two datasets to measure the efficiency of TTPXHunter and the current literature: 1)1)1 ) The augmented Sentence-TTP dataset and 2)2)2 ) The Report-TTP dataset.
@@ V-A1 Augmented Sentence-TTP Dataset
To prepare the sentence-TTP dataset, we consider the TTPHunter dataset as a base dataset, prepared using MITRE ATT&CK knowledgebase [2]. The base dataset consists of two columns: sentences and their corresponding TTP ID. The dataset consists of $10,906$ sentences over $193$ TTPs. We use our proposed data augmentation algorithm, explained in Algorithm 1, and extend the dataset to $39,296$ sentences distributed over $193$ TTP classes. The TTP ID $T1059$ (Command and Scripting Interpreter) consists of the highest number of sentences as $800$, and TTP ID $T1127$ (Trusted Developer Utilities Proxy Execution) consists of the lowest number of samples as $3$. On average, the number of samples in our dataset is $203$. The distribution of data samples is present in Appendix -A and a glimpse of the dataset is shown at step 5⃝ in Fig 2.
@@ V-A2 Report-TTP Dataset
Generally, threat data is present in the form of threat analysis reports. Therefore, We evaluate TTPXHunter on a document dataset, which demonstrates the performance of TTPXHunter in filtering irrelevant sentences from threat reports. It also tells us how efficiently TTPXHunter can extract threat intelligence from threat reports. We manually collect $149$ threat reports published by various prominent security firms, and we manually label the set of TTPs present in each report to prepare the ground truth. The document-TTP dataset contains two columns: threat report and list of TTP present in the corresponding threat report.
@@ V-B Evaluation
We evaluate TTPXHunter using various performance metrics and compare its performance with several current studies. As our dataset is imbalanced, shown in Appendix Fig. 9, we consider macro-averaged precision, recall, and f1-score [11]. This approach ensures balanced evaluation across all classes [29]. By giving equal importance to each TTP class, these metrics prevent the majority class’s dominance from overshadowing the minority class’s performance. It promotes the development of effective and fair models across different TTP classes and is essential for nuanced TTP classification tasks.
@@ V-B1 Augmented Sentence-based Evaluation
We finetune TTPXHunter on the prepared augmented dataset and compare the other two BERT-based models, i.e., TRAM [8] and TTPHunter [2] present in the literature. We divide the prepared augmented dataset into train and test sets with an $80:20$ ratio.
TTPXHunter vs TRAM : We fine-tune the TTPXHunter on the train set and evaluate its performance using the chosen performance metrics. Further, we finetune the literature TRAM [8] on the same dataset and evaluate its performance. By employing TRAM on the augmented dataset, we extend the capability of TRAM from the 50 most frequently used TTPs to the full spectrum of TTPs. As a result, it gives common ground for comparing TTPXHunter and TRAM. The result obtained by both methods and their comparison is shown in Fig. 4. As we can see, the TTPXHunter outperforms the TRAM, which reflects the difference between contextual embedding of general scientific BERT and domain-specific BERT embeddings. This result reflects that the domain-specific language model provides a better contextual understanding of embedding than the language model trained on general scientific terms.
@@
TTPXHunter vs TTPHunter : We assess the performance of our proposed TTPXHunter alongside state-of-the-art TTPHunter. Our proposed extraction method, TTPXHunter, performs better than state-of-the-art TTPHunter. TTPXHunter’s superiority is due to using a cyber-domain-specific finetuned language model. Sentences containing domain-specific terms, such as ”window” and ”registry,” introduce a distinct context that differs from general English. This distinction allows our method, based on the domain-specific language model, to capture and interpret the contextual meaning more accurately than traditional models. In addition, TTPXHunter can identify the range of $193$ TTPs with $0.92$ f1-score, whereas TTPHunter is limited to only $50$ TTPs. The improvement in the result and the capability to identify all ranges of TTPs make TTPXHunter superior to TTPHunter.
@@
Further, to understand the efficiency of the data augmentation method and compare the performance of TTPXHunter and TTPHunter on the same base, we evaluate TTPXHunter on the ground of TTPHunter. We only selected 50 TTP sets for which we developed TTPHunter and evaluated both. The obtained result is shown in Fig. 5. As we can see, TTPXHunter outperforms TTPHunter on the 50-TTP set ground of TTPHunter. It reflects that augmenting more samples for a 50-TTP set and employing a domain-specific language model enables the classifier to better understand the context of TTP.
@@ V-B2 Report-based Evaluation
In the real world, we have threat reports in the form of natural language rather than sentence-wise datasets, and these reports contain information along with TTP-related sentences. So, we evaluate the TTPXHunter on the report dataset, which contains each sample as threat report sentences and a list of TTP explained in the report. Extracting TTPs from threat reports is a multi-label problem because a list of TTP classes is expected as output for any given sample, i.e., threat report in this case. The evaluation of such classification also requires careful consideration because of multi-label classification.
Evaluation Metrics: In the multi-label problem, the prediction vector appears as a multi-hot vector rather than a one-hot vector in a multi-class problem. In the multi-label case, there may be a situation where not all expected TTP classes were predicted; instead, a subset of them is correctly predicted. However, the prediction may be wrong because the whole multi-hot vector does not match. For example, If the true label set contains $\{T1,T2,T3\}$ and the predicted label is $\{T2,T3\}$, then it may be considered to be a mismatch even though $T2$ and $T3$ are correctly classified. So, relying on accuracy may not be a good choice for multi-label problems [28]; instead, we consider hamming loss as a performance metric to deal with such a scenario. The hamming loss measures the error rate label-wise [19, 21, 28]. It calculates the ratio of incorrect labels to all labels. For given $k$ threat reports, the hamming loss is defined as:
@@
Further, we also evaluate macro precision, recall, and f1-score by leveraging a multi-label confusion matrix package from sklearn [11]. Then, we calculate true positive, false positive, and false negative for each class and calculate these performance metrics. Further, we calculate the macro average between all classes to get macro-averaged performance metrics for all chosen measures, i.e., precision, recall, and f1-score. We prefer the macro-average method to ignore biases towards the majority class and provide equal weight to all classes.
We consider four state-of-the-art methods, i.e., [10, 6, 4, 8] for comparison against TTPXHunter based on these metrics over the report dataset. This comparison aims to understand the effectiveness of TTPXHunter over state-of-the-art for TTP extraction from finished threat reports. These methods provide the list of TTPs extracted from the given threat report and the model confidence score for each. Out of all extracted TTPs, only relevant TTPs are selected based on the threshold mechanism decided by each method. We evaluate the state-of-the-art method’s performance based on the threshold value given in their respective papers. For TTPXHunter, we obtain the same threshold experimentally chosen for TTPHunter, i.e., $0.644$. The results obtained from all implemented methods on our report dataset are shown in Table IV. It demonstrates that TTPXHunter outperforms all implemented methods across all chosen metrics, i.e., the lowest hamming loss and the highest other performance metrics. It achieves the highest f1-score of $97.07\%$, whereas out of all state-of-the-art methods, LADDER [4] performs better than other state-of-the-art methods and achieves $2^{nd}$ highest performance of $93.09\%$ F1-score. This performance gain over state-of-the-art methods demonstrates the efficiency of the TTPXHunter, and we plan to make it open for the benefit of the community.
As this experiment involves $193$ target TTP classes, it is challenging to visualize the class-wise performance of the employed models. Therefore, we follow a different way to assess the class-wise efficiency of the employed models. We count the number of TTP classes whose chosen performance metrics lie within a range. We employ a range interval of $0.1$, i.e., $10\%$, to calculate the number of TTP classes whose score falls into the range.
We perform this calculation across all five methods and three chosen performance metrics. The obtained results are present in Figures 6, 7, and 8. The observation reveals that most TTP classes analyzed by rcATT fall within the $0-0.10$ range, contributing to its overall lower performance. This performance is due to the reliance on the TF-IDF method to transform sentences into vectors. TF-IDF is a numerical statistic that reflects how important a word is to a document in a collection or corpus, balancing its frequent appearance within a document against its commonness across all documents [26, 20, 12]. This method lacks the ability to understand the context and semantic relationships between words, making it unable to grasp the overall meaning of sentences [16, 33]. ATTACKG, conversely, exhibits a lower score within the $0-0.10$ range for certain TTP classes, which adversely affects its overall performance. However, TRAM shows a minimum performance range of $0.3-0.4$ for TTP classes, which indicates better performance than rcATT and ATTACKG. LADDER ensures a performance score of at least $0.4-0.5$ for a TTP class, which positions it ahead of the aforementioned methods, including rcATT, ATTACKG, and TRAM. Our proposed model, TTPXHunter, assures a minimum score of $0.6-0.7$ for TTP classes, with the majority exhibiting scores between $0.9-1.0$, which underscores TTPXHunter’s significant advantage over the other methods. It demonstrates the effectiveness of domain-specific models for domain-specific downstream tasks.
Figure 6: Class-wise Precision Score Comparison with State-of-the-art Methods
Figure 7: Class-wise Recall Score Comparison with State-of-the-art Methods
Figure 8: Class-wise F1 Score Comparison with State-of-the-art Methods
@@ VI Limitations & Future Directions
In addition to the advantages of TTPXHunter, like threat intelligence extraction, threat profiling, and sharing, it has some limitations that one should be careful about while using for actionable insights. The MITRE ATT&CK framework is not a kind of fixed knowledge base. Instead, MITRE threat researchers are continuously updating it. One may need to finetune the model again if new TTPs are added to get the newer TTPs predicted. While finetuning on newer TTPs may require using our proposed data augmentation method to create new augmented sentences specific to newer TTPs. Therefore, we plan to make it public so one can adapt our method for any new upcoming versions of the ATT&CK framework.
Further, the TTPXHunter contains a one-to-one classifier model, which maps a given sentence to a single TTP. Like a single sentence, a sentence containing one-to-many mapping may explain more than one TTP. For example, the sentence is "The attacker gained initial access through a phishing email and obtained persistence via run registry modification". TTPXHunter maps this sentence to T1566 (Phishing) or T1037 (Boot or Logon Initialization Scripts) in such a scenario. Extending the capability of TTPXHunter to identify such one-to-many mapping can help us improve the model’s performance. We plan to take up this challenge to improve the efficiency of TTPXHunter in the future.